OpenStudy (anonymous):

xyz+xy+xz+zy+x+y+z=1000

OpenStudy (anonymous):

that's a tough puzzle

OpenStudy (unklerhaukus):

factorise

OpenStudy (anonymous):

Explain?

OpenStudy (unklerhaukus):

xyz+xy+xz+zy+x+y+z=1000 xy(z+1)+z(x+y)+(x+y)+z=1000 ...

OpenStudy (anonymous):

Where can you go from there?

OpenStudy (anonymous):

xyz+xy+xz+zy+x+y+z=1000 (xy)(z+1)+z(x+y)+(x+y)+z (xy)(z+1)+(x+y)(z+1)+z (z+1)(xy+x+y)+z (z+1)(x(y+1)+y)+z

OpenStudy (anonymous):

But what are the varibles equal to?

OpenStudy (unklerhaukus):

hmm,

OpenStudy (anonymous):

Bonus question for a college algebra test that the romanian professor says is 4th grade level for him.

OpenStudy (anonymous):

xyz+xy+xz+zy+x+y+z=1000 (xy)(z+1)+z(x+y)+(x+y)+z (xy)(z+1)+(x+y)(z+1)+z (z+1)(xy+x+y)+z (z+1)((x+1)(y+1)-1)+z (z+1)((x+1)(y+1))-(z+1)+z =1000 (z+1)((x+1)(y+1))-1 =1000 (z+1)((x+1)(y+1))=999 we got it?

OpenStudy (anonymous):

Reward me right now!!! i can't wait

OpenStudy (anonymous):

Thanks

OpenStudy (unklerhaukus):

if you need a unique solution , factories the right hands sides into any three factors,and compare the left and hight hand sides, (there are many possibilities)

OpenStudy (unklerhaukus):

*factorise